Question 9 Question 6 Define opportunity cost In marginal analysis, if possible, the optimal outcome is found when Select the correct answer below: Select the correct answer below O All possible consumption combinations of goods that someone can afford when all income is spent O A cost that is made in the past and cannot be recovered, O All possible combinations of consumption that someone can afford given the prices of goods and the individuals income. O The measure of cost by what is given up in exchange for what is obtained. Question 10 О тв-тс When considering the characteristics of the budget constraint, which of the following statements is true? Question 7 Select the correct answer below The principle of diminishing marginal utility states that. O It is a legal limit on how much individuals may spend without going into bankruptcy O The budget constraint slopes upward. O The budget constraint is a straight line. O The budget constraint is only for countries, not individuals Select the correct answer below people dont enjoy consuming more of a good O as a person receives more of a good, the additional utility from each additional unit of the good increases O people enjoy consuming more of a good O as a person receives more of a good, the additional utility from each additional unit of the good declines Question8 If a family spends its entire budget in a given time frame, the family can afford either 10 restaurant meals or 30 home meals. Assuming the family spends its entire budget on just these two goods, what is the opportunity cost of one extra restaurant meal in the time frame? (Round your answer to two decimal places.) Provide your answer below home meels
Answer #1

Answer:

6)

Option C: MB=MC

Since Net benefit = TB-TC

Since we need to maximize the net benefit so differentiation of net benefit should be zero and differentiation of TB is MB and TC is MC

So MB-MC=0

MB=MC

7)

Option A: People don't enjoy consuming more goods

8)

Opportunity cost for restaurant meal =30/10=3 home meals

9) Option D: The measure of cost by what is given up in exchange for what is obtained.